OpenStudy (anonymous):

To the nearest tenth of a cubic inch, how much volume is added to a spherical balloon whose radius grows from 10 to 20 inches long?

OpenStudy (anonymous):

V=4/3π r^3 if r = 10 V =4000π/3 if r = 20 V= 32000π/3 change in volume is 28000π/3